OpenStudy (anonymous):

Solve for x: x/3 - 7 = 10

OpenStudy (saifoo.khan):

x = 17*3 x = 51

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

x/3 - 7 = 10 x/3 - 7 + 7 = 10 + 7 x/3 * 3 = 17*3 x = 51